OpenStudy (lxelle):

The reflex angle x is such that cos x =k, where 0

OpenStudy (lxelle):

\[\cos ^{2}x + \sin ^{2}x=1\] \[\sin ^{2}x= 1 - \cos ^{2}x\] Therefore, \[sinx = -\sqrt{1-\cos ^{2}x}\]

OpenStudy (lxelle):

Why is it negative square root?

OpenStudy (dan815):

+ or - but why arent u finding sinx in terms of k

OpenStudy (lxelle):

same thing lol, but why is it - squareroot?

OpenStudy (lxelle):

I know it's something because of 0<k<1, but what is that?

OpenStudy (paxpolaris):

a reflex angle is between 180 and 360. and since cosine is positive ... the angle is in the 4th quandrant

OpenStudy (paxpolaris):

sin of such an angle will be negative

OpenStudy (lxelle):

Thanks.

OpenStudy (dan815):

i see